REVIEW

41: What is the equation of the line with slope 2/3 and passes through the midpoint of the segment joining points (-1, 6) and (-7, -4)?

Answer by checkley75(3666)

THE MIDPOINT OF THE LINE THROUGH (-1,6) & (-7,-4) IS
(7-1)/2=6/2=3 FOR THE X POINT
(6+4)/2=10/2=5 FOR THE Y POINT
NOW WE USE THESE 2 POINTS AND THE GIVEN SLOPE OF 2/3 & SOLVE FOR THE Y INTERCEPT(b)
5=2/3*3+b
b=5-6/3
b=(15-6)/3
b=9/3
b=3 THUS THE EQUATION IS
Y=2X/3+3